OpenStudy (anonymous):

what is the slope of a line passing through (2,4) and (4,9)

OpenStudy (anonymous):

y2-y1/x2-x1 where x1= 2, y1=4 x2= 4 y2=9

OpenStudy (anonymous):

so 9-4/4-2=5/2

OpenStudy (anonymous):

\[\frac{ 9-4 }{ 4-2}\]

OpenStudy (anonymous):

exactly @ahaal25

OpenStudy (anonymous):

which simplified would give you a slope of 1.5

OpenStudy (anonymous):

2.5**

OpenStudy (anonymous):

or leave it as the exact answer of 5/2, whatever the book/professor prefers

OpenStudy (anonymous):

mhmm